# F-Actor: Controllable Conversational Behaviour in Full-Duplex Models

### v1

@@ 1 Introduction
Recent work has begun to address full-duplex modelling using a variety of architectural choices: training special predictors for overlapping speech (Ruede et al., 2017; Chen et al., 2025b), modelling user and system speech as separate streams (Défossez et al., 2024; Hu et al., 2025) or by interleaving user and system chunks in a single sequence (Veluri et al., 2024; Lee et al., 2025). While the latter approaches enable models to handle interruptions and backchannels, they typically do not model them explicitly on the system side. That is, models are trained to robustly handle overlapping speech produced by the user, but are rarely studied in terms of whether the system itself interrupts or backchannels. As a result, existing evaluations focus on reactive rather than proactive conversational behaviour (Peng et al., 2025; Lin et al., 2025).
@@ 2 Background
Full-Duplex Modelling. Full-duplex interaction, simultaneous speaking and listening, is often modelled using parallel user and system speech streams (Hu et al., 2025; Défossez et al., 2024). To enable the backbone LLMs to process speech streams, raw audio is typically mapped to discrete acoustic units (DAUs), which can be treated as tokens in the model vocabulary. Each speech stream then emits DAUs at every timestep, including silence DAUs, and embeddings for each stream are combined by summation (Défossez et al., 2024; Hu et al., 2025), or by token-wise fusion (Wang et al., 2025) to allow concurrent processing. To produce speech, the generated DAUs are mapped back into speech waveforms with pre-trained vocoders.
@@
Discrete Acoustic Units. DAUs are typically produced by neural audio codecs such as Mimi (Défossez et al., 2024), EnCodec (Défossez et al., 2022), or SoundStream (Zeghidour et al., 2021). These codecs compress audio into discrete codes optimized for reconstruction and are widely used as both inputs and outputs in recent speech language models (Défossez et al., 2024; Hu et al., 2025).
The codec models represent each audio frame using multiple discrete codes, drawn from multiple codebooks. The choice of quantization determines how these codes are modelled. Dependent codebooks, typically produced by residual vector quantization (Zeghidour et al., 2021, RVQ), require hierarchical prediction: the language model predicts the first code, while additional modules generate the remaining layers (Défossez et al., 2024; Wang et al., 2023). In this work, we choose independent codebooks from finite scalar quantization (Mentzer et al., 2024, FSQ), which simplify modelling by allowing all codes for a frame to be predicted jointly (Hu et al., 2025). Additional details on encoding and reconstruction are provided in App.˜A.
@@ 3 Related Work
End-to-end architectures typically support these behaviours natively and have therefore become increasingly popular. Among them, Moshi (Défossez et al., 2024) and its variants (Ohashi et al., 2025; Shi et al., 2025) represent a common design choice: employing separate system, user, and text streams, and codec-based models for speech input and output (Défossez et al., 2024, 2022; Zeghidour et al., 2021). Other approaches use continuous features at the input (Hu et al., 2025; Fu et al., 2025), or at both input and output (Yu et al., 2024). Some architectures avoid multiple input–output streams by interleaving modalities (Veluri et al., 2024; Lee et al., 2025; Yu et al., 2024), or by embedding them jointly at the token level (Wang et al., 2025). For more details, we refer the reader to Arora et al. (2025a).
@@ 4 Model
F-Actor, our full-duplex model, is based around an LLM backbone, augmented with a speaker embedding model and an audio encoder. Unlike previous work (Défossez et al., 2024; Hu et al., 2025), F-Actor is designed with instruction following in mind, that is, incorporating both a speaker embedding and an instruction-following prompt. In the following, we describe each component of the system in detail. An overview is provided in Fig.˜1.
Architecture. We use Llama3.2-1B-Instruct333[Uncaptioned image] meta-llama/Llama-3.2-1B-Instruct (Grattafiori et al., 2024) as our backbone LLM. For speech encoding and decoding, we employ Nemo-nano-codec-22khz-0.6kbps-12.5fps444[Uncaptioned image] nvidia/nemo-nano-codec-22khz0.6kbps-12.5fps (Casanova et al., 2025). The codec consists of four independent codebooks of size 4032, producing four DAU streams per input audio stream, operating at a frame rate of 12.5 fps. In contrast to Hu et al. (2025), we keep the NanoCodec encoder frozen and train only the LLM.
NanoCodec has two advantages: (i) it uses FSQ quantization (Mentzer et al., 2024) so the four codebooks are independent, enabling parallel prediction without requiring a depth-transformer architecture (Défossez et al., 2024), and (ii) it has shown strong performance on speech synthesis (Hu et al., 2025).
@@
Text Stream. Following Hu et al. (2025); Défossez et al. (2024), we not only generate audio, but also experiment with generating the corresponding text, using an additional system-side text stream. To do so, we need to align the text tokens and audio DAUs (alignment strategies are described in Section˜6). After embedding the text stream, we add it to the system’s speech stream. The LLM uses this summed embedding as input. We then use the LLM’s original language-modeling head to predict the next text token.
@@ 5 Evaluation
We evaluate instruction-following performance along four dimensions: (1) Speaker Initiation: Accuracy in initiating the conversation according to the prompt. (2) Speaker Embedding Consistency: Cosine similarity between the target speaker embedding and the generated speech, averaged across speakers and conversations. To detect potential drift, we also compute the distance (1–cosine similarity) between the first and last segments for each speaker. (3) Narrative Adherence: An LLM judge (Llama-3.1-8B-Instruct555[Uncaptioned image] meta-llama/Llama-3.1-8B-Instruct (Grattafiori et al., 2024) ) evaluates the alignment (of the transcript) with the narrative specified in the prompt (prompts in Fig.˜4). (4) Backchannels (BCs) and Interruptions: We measure the correlation between prompt-specified and generated counts using Pearson’s r, with two-sided p-values computed from the exact distribution.
For (4), we experiment with different algorithms for detecting BCs and interruptions. We initially adopted the FD-Bench (Peng et al., 2025) implementation, which uses Silero-VAD (Team, 2024), but default thresholds did not generalize well to the dataset used in this work (Lee et al., 2025, details in Section˜6), often misclassifying interruptions or BCs. Detecting these events is challenging because short BCs and overlapping speech can be easily missed or misaligned with transcripts. We compare timestamps from Silero-VAD, Parakeet666[Uncaptioned image] nvidia/parakeet-tdt-0.6b-v2, and Vosk777[Uncaptioned image]vosk-model-en-us-0.22, performing a grid search over interruption and overlap threshold parameters. Parakeet achieves the most reliable performance and is used for all subsequent evaluations. Detailed results for all alignment methods are reported in Tab.˜4, with grid-search details provided in Section˜C.2.
@@ 6.1 Data
Text-Speech Alignment. As described in Section˜4, we experiment with generating not only speech, but also the corresponding text. We explore both word- and utterance-level alignment of speech and text, as prior work has shown mixed results (Défossez et al., 2024; Hu et al., 2025). To this end, we align text and speech tokens using forced alignment with Kaldi (Povey et al., 2011) and a public model888[Uncaptioned image] kaldi-model-m13 trained on Librispeech Panayotov et al. (2015). We exclude conversations with text–speech mismatches due to occasional TTS deviations from the transcript, resulting in 74.4% of the original training set.
@@
Emb. Stream Align. Delay DSU $\downarrow$ Text $\downarrow$ $\uparrow$ audio/text $\downarrow$ Diff (s) $\downarrow$
@@ 6.2 Training Details
To maximize efficiency, our approach utilizes a single training stage that avoids additional speech or text pretraining as required by prior models such as Défossez et al. (2024). Only the LLM backbone and the linear DSU heads are trained, while the speech encoder remains frozen. Training is performed on four NVIDIA A100-SXM4-40GB GPUs with early stopping based on validation performance, resulting in around 48 hours of training per model. We use a maximum sequence length of 2,048 tokens, a per-GPU batch size of 1, and 8 gradient accumulation steps. Training and inference parameters are provided in Tab.˜5 in App.˜D.
@@ 6.3 Experiments
Text-Stream. We examine the effect of adding a text stream on the system side (see Section˜4), enabling the model to predict not only audio but also the corresponding text (Défossez et al., 2024; Hu et al., 2025). We further experiment with introducing special tokens that precede BCs and interruptions (BC/I tok.) to help the model generate the numbers specified in the prompt.
[…]
Behaviour-SD Dialogues Testset 100.00 0.62 0.62 4.04 0.92 0.74
Lower Baseline 50.0a 0.35b 0.75c 1.26d - -
s/u n/a n/a 154.00 0.52 0.66 1.94 0.56 0.21
s/u utt. 0 188.76 0.52 0.67 2.12 0.31 0.13
s/u utt. 1 199.00 0.54 0.63 2.20 0.40 0.22
s/u utt. 2 199.60 0.53 0.61 2.23 0.36 0.16
s/u word 0 189.80 0.52 0.66 2.00 0.53 0.23
s/u word 1 199.60 0.54 0.63 2.15 0.54 0.28
s/u word 2 199.80 0.54 0.61 2.39 0.54 0.25
s/u word 2 199.40 0.36 0.64 2.20 0.50 0.19
s word 2 199.60 0.54 0.64 2.27 0.57 0.17
s/u (BC/I tok.) word 2 100.00 0.54 0.62 2.36 0.48 0.28
* All correlations are statistically significant with $p<0.01$
Table 2: Instruction-following results for full-duplex models trained to predict both system and user (s/u) or only the system role (s). Models are compared with and without a text stream, using word- or utterance-level alignment, and with different audio delays relative to the text. BC/I tok indicate special tokens for backchannels and interruptions.
@@
Audio Delay. We experiment with delaying the audio stream by up to two tokens relative to the text stream, generating text tokens first while padding the audio stream before audio token generation begins (Défossez et al., 2024; Hu et al., 2025).
RVQ vs. FSQ. All primary experiments use an FSQ encoder to benefit from independent codebooks. To test the generality of our multi-head DSU prediction architecture, we also evaluate an RVQ-based setup using Mimi101010[Uncaptioned image] kyutai/mimi (Défossez et al., 2024) with 8 codebooks. This allows us to verify whether predicting DSUs for each codebook in parallel remains effective with RVQ quantization.
@@ 7.3 Comparing Turn-Taking to SOTA Models
In Tab.˜3, we compare turn-taking behaviour to Moshi (Défossez et al., 2024) and dGSLM (Nguyen et al., 2023), with all metrics reported as cumulative durations per category. For all four metrics, Inter-Pausal Units (IPUs), intra-turn Pause, between-turn Gap and Overlap, our model exhibits behavior similar to the training data and the baseline models. Like Moshi and non-cascaded dGLSM, our model’s pauses are generally longer than the gaps, reflecting real-world conversational statistics (Heldner and Edlund, 2010).
@@ 7.3.1 The Role of Sampling Parameters
Similar to Défossez et al. (2024), we observe that sampling temperature has a significant impact on model behaviour. Results for different temperatures (0.60.61.01.0) using our best-performing model (word-level alignment; audio delay of 2) on the dev set are reported in Tab.˜6 in App.˜E. As temperature increases, the dialogue becomes more dynamic, with more BCs and interruptions, more balanced speaking time between the speakers, and improved coherence and adherence to the narrative. However, higher temperatures also lead to slightly lower UTMOS scores and higher WER, reflecting a mild impact on speech quality. In our setup, a temperature of $0.9$ provides the best trade-off, and all results reported in this paper use this setting.
@@ 7.4 RVQ vs FSQ
Finally, we train our best-performing configuration (word-level alignment, audio delay of 2, s/u loss) using an RVQ encoder. Results are reported in Tab.˜7 in App.˜E. Although we do not explicitly model codebook dependencies and instead predict all codebooks in parallel, performance is comparable to that of the FSQ-based models. However, as expected, speech quality degrades substantially, with UTMOS dropping to 2.5 and similarly low scores observed for speaker embedding metrics.

### v2

@@ 1 Introduction
Recent work has begun to address full-duplex modelling using a variety of architectural choices: training special predictors for overlapping speech (Ruede et al., 2017; Chen et al., 2025b), modelling user and system speech as separate streams (défossez2024moshispeechtextfoundationmodel; Hu et al., 2025) or by interleaving user and system chunks in a single sequence (Veluri et al., 2024; Lee et al., 2025). While the latter approaches enable models to handle interruptions and backchannels, they typically do not model them explicitly on the system side. That is, models are trained to robustly handle overlapping speech produced by the user, but are rarely studied in terms of whether the system itself interrupts or backchannels. As a result, existing evaluations focus on reactive rather than proactive conversational behaviour (Peng et al., 2025; Lin et al., 2025).
@@ 2 Background
Full-Duplex Modelling. Full-duplex interaction, simultaneous speaking and listening, is often modelled using parallel user and system speech streams (Hu et al., 2025; défossez2024moshispeechtextfoundationmodel). To enable the backbone LLMs to process speech streams, raw audio is typically mapped to discrete acoustic units (DAUs), which can be treated as tokens in the model vocabulary. Each speech stream then emits DAUs at every timestep, including silence DAUs, and embeddings for each stream are combined by summation (défossez2024moshispeechtextfoundationmodel; Hu et al., 2025), or by token-wise fusion (Wang et al., 2025) to allow concurrent processing. To produce speech, the generated DAUs are mapped back into speech waveforms with pre-trained vocoders.
@@
Discrete Acoustic Units. DAUs are typically produced by neural audio codecs such as Mimi (défossez2024moshispeechtextfoundationmodel), EnCodec (Défossez et al., 2022), or SoundStream (Zeghidour et al., 2021). These codecs compress audio into discrete codes optimized for reconstruction and are widely used as both inputs and outputs in recent speech language models (défossez2024moshispeechtextfoundationmodel; Hu et al., 2025).
The codec models represent each audio frame using multiple discrete codes, drawn from multiple codebooks. The choice of quantization determines how these codes are modelled. Dependent codebooks, typically produced by residual vector quantization (Zeghidour et al., 2021, RVQ), require hierarchical prediction: the language model predicts the first code, while additional modules generate the remaining layers (défossez2024moshispeechtextfoundationmodel; Wang et al., 2023). In this work, we choose independent codebooks from finite scalar quantization (Mentzer et al., 2024, FSQ), which simplify modelling by allowing all codes for a frame to be predicted jointly (Hu et al., 2025). Additional details on encoding and reconstruction are provided in App.˜A.
@@ 3 Related Work
End-to-end architectures typically support these behaviours natively and have therefore become increasingly popular. Among them, Moshi (défossez2024moshispeechtextfoundationmodel) and its variants (Ohashi et al., 2025; Shi et al., 2025) represent a common design choice: employing separate system, user, and text streams, and codec-based models for speech input and output (défossez2024moshispeechtextfoundationmodel; Défossez et al., 2022; Zeghidour et al., 2021). Other approaches use continuous features at the input (Hu et al., 2025; Fu et al., 2025), or at both input and output (Yu et al., 2024). Some architectures avoid multiple input–output streams by interleaving modalities (Veluri et al., 2024; Lee et al., 2025; Yu et al., 2024), or by embedding them jointly at the token level (Wang et al., 2025). For more details, we refer the reader to Arora et al. (2025a).
@@ 4 Model
F-Actor, our full-duplex model, is based around an LLM backbone, augmented with a speaker embedding model and an audio encoder. Unlike previous work (défossez2024moshispeechtextfoundationmodel; Hu et al., 2025), F-Actor is designed with instruction following in mind, that is, incorporating both a speaker embedding and an instruction-following prompt. In the following, we describe each component of the system in detail. An overview is provided in Fig.˜1.
Architecture. We use Llama3.2-1B-Instruct333[Uncaptioned image] meta-llama/Llama-3.2-1B-Instruct (Grattafiori et al., 2024) as our backbone LLM. For speech encoding and decoding, we employ Nemo-nano-codec-22khz-0.6kbps-12.5fps444[Uncaptioned image] nvidia/nemo-nano-codec-22khz0.6kbps-12.5fps (Casanova et al., 2025). The codec consists of four independent codebooks of size 4032, producing four DAU streams per input audio stream, operating at a frame rate of 12.5 fps. In contrast to Hu et al. (2025), we keep the NanoCodec encoder frozen for efficiency and train only the LLM.
NanoCodec has two advantages: (i) it uses FSQ quantization (Mentzer et al., 2024) so the four codebooks are independent, enabling parallel prediction without requiring a depth-transformer architecture (défossez2024moshispeechtextfoundationmodel), and (ii) it has shown strong performance on speech synthesis (Hu et al., 2025).
@@
For (c), we specify backchanneling and interruption frequency as exact counts rather than proportions. This choice enables controlled evaluation across full-duplex models (e.g., assessing how well a model responds when interrupted exactly 5 times and backchanneled 4 times), while remaining flexible, as counts and proportions are trivially interconvertible.
@@
Text Stream. Following Hu et al. (2025); défossez2024moshispeechtextfoundationmodel, we not only generate audio, but also experiment with generating the corresponding text, using an additional system-side text stream. To do so, we need to align the text tokens and audio DAUs (alignment strategies are described in Section˜6). After embedding the text stream, we add it to the system’s speech stream. The LLM uses this summed embedding as input. We then use the LLM’s original language-modeling head to predict the next text token.
@@ 5 Evaluation
We evaluate instruction-following performance along four dimensions: (1) Speaker Initiation: Accuracy in initiating the conversation according to the prompt. (2) Speaker Embedding Consistency: Cosine similarity between the target speaker embedding and the generated speech, averaged across speakers and conversations. To detect potential drift, we also compute the distance (1–cosine similarity) between the first and last segments for each speaker. (3) Narrative Adherence: An LLM judge (Llama-3.1-8B-Instruct555[Uncaptioned image] meta-llama/Llama-3.1-8B-Instruct (Grattafiori et al., 2024) ) evaluates the alignment (of the transcript) with the narrative specified in the prompt (prompts in Fig.˜4). We run the judge with three different seeds and report the mean and standard deviation across these. We also run a human evaluation to confirm the reliability of the judge. Details can be found in Section˜7.2. (4) Backchannels (BCs) and Interruptions: We measure the correlation between prompt-specified and generated counts using Pearson’s r, with two-sided p-values computed from the exact distribution.
For (4), we experiment with different algorithms for detecting BCs and interruptions. We initially adopted the FD-Bench (Peng et al., 2025) implementation, which uses Silero-VAD (Team, 2024), but default thresholds did not generalize well to the dataset used in this work (Lee et al., 2025, details in Section˜6), often misclassifying interruptions or BCs. Detecting these events is challenging because short BCs and overlapping speech can be easily missed or misaligned with transcripts. We compare timestamps from Silero-VAD, Parakeet666[Uncaptioned image] nvidia/parakeet-tdt-0.6b-v2, and Vosk777[Uncaptioned image]vosk-model-en-us-0.22, performing a grid search over interruption and overlap threshold parameters. Parakeet achieves the most reliable performance and is used for all subsequent evaluations. Detailed results for all alignment methods are reported in Tab.˜5, with grid-search details provided in Section˜C.2.
@@ 6.1 Data
Text-Speech Alignment. As described in Section˜4, we experiment with generating not only speech, but also the corresponding text. We explore both word- and utterance-level alignment of speech and text, as prior work has shown mixed results (défossez2024moshispeechtextfoundationmodel; Hu et al., 2025). To this end, we align text and speech tokens using forced alignment with Kaldi (Povey et al., 2011) and a public model888[Uncaptioned image] kaldi-model-m13 trained on Librispeech Panayotov et al. (2015). We exclude conversations with text–speech mismatches due to occasional TTS deviations from the transcript, resulting in 74.4% of the original training set.
@@
Emb. Stream Align. Delay DAU $\downarrow$ Text $\downarrow$ $\uparrow$ audio/text $\downarrow$ Diff (s) $\downarrow$
@@ 6.2 Training Details
To maximize efficiency, our approach utilizes a single training stage that avoids additional speech or text pretraining as required by prior models such as défossez2024moshispeechtextfoundationmodel. Only the LLM backbone and the linear DAU heads are trained, while the speech encoder remains frozen. Training is performed on four NVIDIA A100-SXM4-40GB GPUs with early stopping based on validation performance, resulting in around 48 hours of training per model. We use a maximum sequence length of 2,048 tokens, a per-GPU batch size of 1, and 8 gradient accumulation steps. Training and inference parameters are provided in Tab.˜6 in App.˜D.
@@ 6.3 Experiments
Text-Stream. We examine the effect of adding a text stream on the system side (see Section˜4), enabling the model to predict not only audio but also the corresponding text (défossez2024moshispeechtextfoundationmodel; Hu et al., 2025). We further experiment with introducing special tokens that precede BCs and interruptions (BC/I tok.) to help the model generate the numbers specified in the prompt.
@@
Audio Delay. We experiment with delaying the audio stream by up to two tokens relative to the text stream, generating text tokens first while padding the audio stream before audio token generation begins (défossez2024moshispeechtextfoundationmodel; Hu et al., 2025).
RVQ vs. FSQ. All primary experiments use an FSQ encoder to benefit from independent codebooks. To test the generality of our multi-head DAU prediction architecture, we also evaluate an RVQ-based setup using Mimi101010[Uncaptioned image] kyutai/mimi (défossez2024moshispeechtextfoundationmodel) with 8 codebooks. This allows us to verify whether predicting DAUs for each codebook in parallel remains effective with RVQ quantization.
@@ 7.1 General Modeling Abilities
[…]
Behaviour-SD Dialogues Testset 100.00 0.62 0.62 3.90±0.03 0.92 0.74
Lower Baseline 50.0a 0.35b 0.75c 1.26±0.02d - -
s/u nan - 54.0 0.52 0.66 1.50±0.03 0.56 0.21
s/u utt. 0 88.76 0.52 0.67 2.12±0.01 0.31 0.13
s/u utt. 1 99.0 0.54 0.63 2.24±0.03 0.4 0.22
s/u utt. 2 99.6 0.53 0.61 2.30±0.00 0.36 0.16
s/u word 0 89.8 0.52 0.66 1.74±0.02 0.53 0.23
s/u word 1 99.6 0.54 0.63 2.12±0.03 0.54 0.28
s/u word 2 99.8 0.54 0.61 2.78±0.02 0.54 0.25
s/u word 2 99.4 0.36 0.64 2.53±0.05 0.5 0.19
s word 2 99.6 0.54 0.64 2.46±0.06 0.57 0.17
s/u (BC/I tok.) word 2 100.0 0.54 0.62 2.49±0.02 0.48 0.28
* All correlations are statistically significant with $p<0.01$
Table 2: Instruction-following results for full-duplex models trained to predict both system and user (s/u) or only the system role (s). Models are compared with and without a text stream, using word- or utterance-level alignment, and with different audio delays relative to the text. BC/I tok indicate special tokens for backchannels and interruptions. The narrative was judge by an LLM judge (details in Section˜5) with three different seeds, we report the mean and standard deviation across these.
@@ 7.2 Instruction Following
Model Human Rating LLM Judge
Behaviour-SD Dialogues 4.17 3.88
Model 2 audio delay 3.00 2.68
Model 2 audio delay + BC/I tok. 2.68 2.33
Table 3: Human Evaluation to rate the narrative of the generated dialogues on 80 dialogues. The human rating confirms the ranking of the LLM judge.
To confirm the reliability of the LLM judge, we conduct a human evaluation to analyse its performance. We select our two best-performing models (both using s/u loss, speaker embeddings, word alignment, and an audio delay of 2; one also incorporating BC and interruption tokens) alongside the Behaviour-SD original dialogues. Using the Pearmut platform (Zouhar and Kocmi, 2026), 12 annotators each rated 60 samples (20 dialogues × 3 models), with every dialogue annotated by 3 annotators, yielding 240 annotated instances across 80 unique dialogues. Tab.˜3 shows the results of this evaluation. Specifically, the group-by-system correlation between the LLM judge and human annotators is $\tau=1.00$, confirming that the LLM ranking of the three systems perfectly matches the human ranking. At the item level, the group-by-item correlation is $\tau=0.507$, indicating moderate but solid agreement on individual dialogues. Similarly, human annotators showed substantial agreement with each other, with a mean Kendall’s $\tau$ of 0.517 at the item level (across 240 reviewer pairs) and a $\tau=1.00$ at the system level (across 3 reviewer pairs), meaning all annotators ranked the three systems in the same order. More details on the human evaluation can be found in Section˜C.3.
@@ 7.3 Comparing Turn-Taking to SOTA Models
In Tab.˜4, we compare turn-taking behaviour to Moshi (défossez2024moshispeechtextfoundationmodel) and dGSLM (Nguyen et al., 2023), with all metrics reported as cumulative durations per category. For all four metrics, Inter-Pausal Units (IPUs), intra-turn Pause, between-turn Gap and Overlap, our model exhibits behavior similar to the training data and the baseline models. Like Moshi and non-cascaded dGLSM, our model’s pauses are generally longer than the gaps, reflecting real-world conversational statistics (Heldner and Edlund, 2010).
@@ 7.3.1 The Role of Sampling Parameters
Similar to défossez2024moshispeechtextfoundationmodel, we observe that sampling temperature has a significant impact on model behaviour. Results for different temperatures (0.60.61.01.0) using our best-performing model (word-level alignment; audio delay of 2) on the dev set are reported in Tab.˜7 in App.˜E. As temperature increases, the dialogue becomes more dynamic, with more BCs and interruptions, more balanced speaking time between the speakers, and improved coherence and adherence to the narrative. However, higher temperatures also lead to slightly lower UTMOS scores and higher WER, reflecting a mild impact on speech quality. In our setup, a temperature of $0.9$ provides the best trade-off, and all results reported in this paper use this setting.
@@ 7.4 RVQ vs FSQ
Finally, we train our best-performing configuration (word-level alignment, audio delay of 2, s/u loss) using an RVQ encoder. Results are reported in Tab.˜8 in App.˜E. Although we do not explicitly model codebook dependencies and instead predict all codebooks in parallel, performance is comparable to that of the FSQ-based models. However, as expected, speech quality degrades substantially, with UTMOS dropping to 2.5 and similarly low scores observed for speaker embedding metrics.